Find the unknown resistance AB when the resistance of the remaining arms are BC=120ohms, CD=110ohms, DA=110ohms, AB=?.
1
Expert's answer
2019-11-01T06:02:13-0400

If we consider a Wheatstone bridge, and the circuit is balanced, i.e. the potential difference between points B and D is 0, then the unknown resistance is can be found from the relation


"\frac{R_{AB}}{R_{BC}}=\frac{R_{AD}}{R_{CD}},\\
\space\\
R_{AB}=R_{BC}\cdot\frac{R_{AD}}{R_{CD}}=120\cdot110/110=120\space \Omega."